Internal Service Fund Financial Statements

At the beginning of fiscal year 2014, the City of Wooster established a central supplies storehouse to service its several funds. The general fund contributed $26,000,000 (nonrefundable) to aid in the establishment of the supplies storehouse. It was agreed that the storehouse would charge other funds for the purchase price of supplies plus 15%. During the year, the storehouse purchased $22,000,000 of supplies, paid operating expenses of $3,000,000, and billed other funds for $21,275,000. All accounts are settled except $1,200,000 remaining to be collected from the general fund for supplies billed.

Required

Prepare the statement of revenues, expenses, and changes in net position for fiscal 2014, and the statement of net position at the end of fiscal 2014 for the central supplies storehouse.

  • Enter your answers in thousands. For example, $1,000,000 is $1,000.
  • Use a negative sign with your answer for operating income (loss) to indicate an operating loss; otherwise, do not use a negative sign with your answers.
  • Enter your answers in thousands. For example, $1,000,000 is $1,000. Use a negative sign with your answer for operating incomeCity of Wooster Central Supplies Storehouse Statement of Net Position At Fiscal Year-End 2014 (in thousands) Assets Current a
Answer #1

Ans :

Net position at begining : 0 (As Investment made during the year)

Cost of purchases : 22,000

Operating Exp. : 3,000

operating Revenue : 21,275

Due from General fund : 1,200

(1) Since it is agreed that storehouse will charge other funds for Purchase price + 15%. Hence

Cost of goods sold : operating revenue *100/(100+15)

Placing figures : 21,275*100/115 = 18,500 (cost of goods sold).....................(1)

(2) Inventory : purchases - Cost of goods sold = 22,000-18,500 = 3,500.....................(2)

(3) cash at fiscal year end : Total inflows during the year - Total outflows duriing the year

Cash Sales : Operating Revenue - Due from general fund

21,275 - 1,200 = 20,075

Total inflows : non refundable general fund (Capital Contributions) + Cash Sales = 26,000 + 20,075 =   46,075

Total outflows : Operating Expenses + Purchases = 3,000 + 22,000 = 25,000

Cash at fiscal year end : 46,075 -25,000 = 21,075..............................................................(3)

(4) Operating Expenses including cost of goods sold ; 3,000 + 18,500 = 21,500...............(4)

City of wooster
Central Supplies Storehouse
Statement of revenues, expenses and changes in net position
For the year 2014
(in thousands) Currency Figure
Operatig Revenue $ 21,275
Operatig Expense (4) $ -21,500
Operating Income $ -225
Capital Contributions $ 26,000
Change in Net positions $ 25,775
Total Net position - beginning $ 0
Total Net position - Ending $ 25,775
City of wooster
Central Supplies Storehouse
Statement of net position
At Fiscal Year-End 2014
(in thousands) Currency Figure
Assets
Current Asstes :
Cash (3) $ 21,075
Due from General Fund $ 1,200
Inventory (2) $ 3,500
Total Assets $ 25,775
Net Position
Unrestricted $ 25,775
Total Net Position $ 25,775